# Constraining the Stellar-to-Halo Mass Relation with Galaxy Clustering and Weak Lensing from DES Year 3 Data

### v2

@@ 3.1 Galaxy-matter cross power spectrum $P_{\rm gm}$
Following closely our previous work in Zacharegkas et al. (2022), which follows Seljak (2000); Mandelbaum et al. (2005), we can write down the expressions for the terms in Equation (10) as follows.
@@
[…]
×$\quad\times u_{\rm dm}(k|M_{h})\left[\langle N_{c}(M_{h})\rangle+\langle N_{s}(M_{h})\rangle u_{\rm sat}(k|M_{h})\right]\;,$ (12)
where $u_{\rm dm/sat}(k|M_{h})$ is the Fourier transform of the dark matter/satellite spatial distribution, for which we assume the NFW profile (Navarro et al., 1996), and the average galaxy number density
[…]
@@
[…]
×$\times\int dM_{h}^{\prime}\;\frac{dn}{dM_{h}^{\prime}}\frac{b_{h}(M_{h}^{\prime})}{n_{g}}$
×$\times\left[\langle N_{c}(M_{h}^{\prime})\rangle+\langle N_{s}(M_{h}^{\prime})\rangle u_{\rm sat}(k|M_{h}^{\prime})\right]\;,$ (15)
where, usually, the first and second term in the brackets are identified as the 2-halo central and 2-halo satellite terms, respectively. In this work we adopt the functions from Tinker et al. (2010) for the halo bias, $b_{h}(M_{h})$, and the Halo Mass Function (HMF), $dn/dM_{h}$, from Tinker et al. (2008). We note here that on the very large scales, the above reduces to
@@
[…]
is the large-scale galaxy bias, and for the linear matter power spectrum $P_{\rm m}^{\rm lin}(k,z)$ we adopt the fitting functions from Eisenstein & Hu (1998) which we modify by adopting the Halofit (Takahashi et al., 2012) approximation to compute the nonlinear matter power spectrum, $P_{\rm m}^{\rm nl}(k,z)$. We furthermore emphasize here that the first integral in Equation (3.1) must approach unity on those linear scales. In practice, we need to choose a galaxy bias model that when combined with the halo-mass function satisfies this requirement, i.e. that the dark matter is unbiased with respect to itself, and thus:
[…]
@@ 3.2 Galaxy auto-correlation power spectrum $P_{\rm gg}$
To model the galaxy clustering signal, we follow van den Bosch et al. (2013).
For the 1-halo components of Equation (11), we follow Seljak (2000) which leads to the 1-halo galaxy auto-correlation power spectrum
[…]
×$\quad\times\left[2\langle N_{c}(M_{h})\rangle+\lambda_{s}(M_{h})u_{\rm sat}(k|M_{h})\right]\;,$ (19)
@@
×$\;\times\int dM_{h}\;\frac{dn}{dM_{h}}\frac{b_{h}(M_{h})}{n_{g}}\left[\langle N_{c}(M_{h})\rangle+\lambda_{s}(M_{h})u_{\rm sat}(k|M_{h})\right]$
×$\;\times\int dM_{h}^{\prime}\;\frac{dn}{dM_{h}^{\prime}}\frac{b_{h}(M_{h}^{\prime})}{n_{g}}\rangle\left[\langle N_{c}(M_{h}^{\prime})\rangle+\lambda_{s}(M_{h}^{\prime})u_{\rm sat}(k|M_{h}^{\prime})\right]$
@@ 3.3 Transition regime
Following Mead et al. (2021), we implement a ”beyond-linear” halo bias model when combining the 1 and 2-halo terms in Sections 3.1 and 3.2. This changes the shape of the 2-halo component as a function of $k$, especially around the scales corresponding to the size of individual dark matter halos. This scale-dependence of the halo bias becomes important on scales $r\lesssim 3\;{\rm Mpc}/h$ and we have found that it is especially important in modeling $\xi_{gg}$, as without it the 2-halo term can dominate even in the quasi-to-non-linear regime for low-mass galaxies. Moreover, this modification is especially useful for fitting our model for $\xi_{\rm gm}$ to the low-stellar mass bins, as without it our fits fail in the small-scale regime. In this paper, we implement the scale-dependent halo bias by utilizing the fitting functions from Tinker et al. (2005), which are calibrated on numerical simulations and were found to be accurate enough to model galaxy clustering for galaxies in the Sloan Digital Sky Survey (SDSS; York et al., 2000) using the HOD framework. In particular, we multiply the 2-halo correlation function by $\zeta(r,z)$, which is defined as
@@
where $\xi_{\rm m}^{\rm nl}(r,z)$ is the non-linear real-space correlation function of $P_{\rm m}^{\rm nl}(k,z)$. This is the same approach also used in Zu & Mandelbaum (2015) and results in the 2-halo galaxy-matter cross- and galaxy auto- correlation function terms
@@
respectively. We find that using this model we are able to obtain a better fit to our data on all scales and improve the fits significantly in the 1-halo to 2-halo regime compared to what we assumed previously in Zacharegkas et al. (2022), which was also followed by Hayashi & White (2008); Zu et al. (2014); Clampitt et al. (2017). As we discuss in Section 6, this model is especially useful in fitting the low-redshift, low-mass galaxy bins.
@@ 3.4 The observables: $\gamma_{t}$ and $w$
[…]
@@ 3.5 Uncertainty in shear calibration, redshift and stellar-mass
To account for uncertainties in the shear calibration, redshift and stellar mass distributions, we make some substitutions to the equations of the previous sections, as we describe in the remaining of this section. Before we proceed, we note that aside from the marginalization of the stellar mass distribution, our approach is the same as that used in Prat et al. (2022).
@@ 3.6 Additional modeling components
Similarly to Zacharegkas et al. (2022), we consider a number of additional modeling components to our data vectors $\gamma_{t}$ and $w$. We briefly summarize each of these components and show the full model in Figure 2.
@@
where we model the 2D projected power spectrum $C_{\rm gm}^{\rm lmag}(\ell)$ following Unruh et al. (2020); Zacharegkas et al. (2022) and where $\alpha_{\rm lmag}$ is a free parameter. More specifically, in our notation, for lens bin $i$ and source bin $j$, we vary one additional parameter due to lens magnification that we call $\alpha_{\rm lmag}^{i,j}$.
Tidal stripping of satellites: As tidal forces in the outskirts of the host halo strips off the dark matter content of the satellite subhalo, the density profile of the subhalos drops off at large scales. Therefore, this term is usually modeled as a truncated NFW profile (see e.g. Mandelbaum et al., 2005; Velander et al., 2013) which is similar to that of the central 1-halo, $\gamma_{t}^{\rm c1h}$, out to the truncation radius $R$ and falls off as $\propto r^{-2}$ at larger radii $r$. For this work, we fix the truncation radius to $R=0.4R_{200c}$. Additionally, since this is a satellite term, it needs to be multiplied by the satellite fraction $F_{\rm sat}$, and we thus write:
@@
Point-mass contribution: This term quantifies the contribution to lensing by the baryonic content of the central galaxy (e.g. Velander et al., 2013). This term is simply modelled as a point-source term:
@@
In this work, $\langle M_{\star}\rangle$ is the average stellar mass content in a given lens bin from Table 1. The final contribution of this term to $\gamma_{t}$ is computed as in Zacharegkas et al. (2022). There is no free parameters in this component.
Density profiles for satellite galaxies: For the spatial distribution of dark matter and satellite galaxies within a dark matter halo we have assumed the NFW proflie (Navarro et al., 1996). Thus, their Fourier transforms have analytic forms with concentration parameters $c_{\rm dm}$ and $c_{\rm sat}$, respectively for dark matter and satellite galaxies. However, the distribution of satellite galaxies is typically less concentrated than that of the dark matter (e.g. Carlberg et al., 1997; Lin et al., 2004; Nagai & Kravtsov, 2005; Hansen et al., 2005). To account for this, in this work we allow $c_{\rm sat}$ to be different than $c_{\rm dm}$ by introducing the free parameter $\kappa_{c}^{i}=c_{\rm sat}/c_{\rm dm}$ per lens bin $i$.
Intrinsic alignment: Galaxies are not randomly oriented even in the absence of lensing. On large scales, galaxies can be stretched in a preferable direction by the tidal field of the large scale structure. On small scales, other effects such as the radial orbit of a galaxy in a cluster can affect their orientation. This phenomenon, where the shape of the galaxies is correlated with the density field, is known as intrinsic alignment (IA, Troxel & Ishak, 2015). The contamination of shear by IA can become important when the source galaxies are physically close to the lenses and gravitational interactions can modify the shape of the galaxies. A good proxy for whether IA’s contribution to $\gamma_{t}$ is significant is the measured boost factors (see Section 5.4 and Appendix E for the relevant discussions and definitions). Therefore, given our bin selection and scale cuts, which are described in Section 5.4, we have ignored the contribution from IA in this work.
4 Data
In this section we introduce the data used in this work. This work is based on the first three years (Y3) of DES data (Sevilla-Noarbe et al., 2021) and inherits all the infrastructure and value-added products associated with that dataset (e.g. Gatti et al., 2021; Myles et al., 2021; Hartley et al., 2022). The most novel aspect in the dataset here is the new stellar mass-selected lens sample we construct and validate in this work. The source sample follows the same selection as described in Zacharegkas et al. (2022); Prat et al. (2022). As such, we describe in more detail the lens sample in Section 4.1 and briefly introduce the source sample in Section 4.2.
4.1 The stellar mass-selected lens sample
To generate our stellar mass-selected sample, we start from the DES Y3 GOLD catalog (Sevilla-Noarbe et al., 2021). We first define an approximately complete galaxy sample by performing the following steps:
  • Select all objects with i-band magnitude sof_cm_mag_corrected_i$<23$
  • Define footprint to only contain regions with limiting magnitude deeper than sof_cm_mag_corrected_i$<23$
[…]
This catalog is then fed into the GalPro (Mucesh et al., 2020) package to generate photometric redshift and stellar mass for each galaxy. We briefly introduce GalPro and our training set in Section 4.1.1, then define the actual sample used for this work in Section 4.1.2. In Section 4.1.3 we describe the procedure in which we obtain weights for this lens sample to ensure it is not significantly impacted by the spatial variation of observing conditions.
@@ 4.1.1 GalPro and the training set
GalPro (Mucesh et al., 2020) is a python-based machine learning code based on the Random Forest algorithm for estimating multivariate posterior probability distributions of galaxy properties. It can operate generically on a wide range of galaxy properties (e.g. redshift, stellar mass, star formation rate, metallicity), while we use it to extract redshift and stellar mass information.
The input quantities that GalPro accepts are luptitudes (Lupton et al., 1999), $\mu$, their uncertainties $\sigma_{\mu}$, as well as lupti-colors and their uncertainties. Luptitudes behave like magnitudes for bright photometry and like fluxes for faint photometry, with the turning point in the behaviour determined by the softening parameter. Thus, by converting to luptitudes, we avoid introducing an additional selection effect by not discarding galaxies with negative fluxes.
@@
For the colors, we calculate three lupti-colors $g-r$, $r-i$ and $i-z$ and their corresponding errors. These are simply calculated as $\mu_{\rho}-\mu_{\nu}$ for two bands $\rho,\nu\in\{g,r,i,z\}$ and $\sqrt{\sigma_{\mu,\rho}^{2}+\sigma_{\mu,\nu}^{2}}$, respectively.
Before we can apply GalPro to the DES data, we need to train the random forest with a training set where we have “truth” information for the mapping between the photometry and the desired quantities (redshift and stellar mass in our case). We choose to train our model on galaxies in the DES Y3 Deep Fields (DF, Hartley et al., 2022), matched to an overlapping subset of galaxies in the Cosmic Evolution Survey COSMOS2020 (COSMOS, Weaver et al., 2022). In matching the two samples, we adjust the stellar mass slightly according to the difference in amplitude between the DF photometry and the COSMOS photometry. We note that this may not address the underlying cause of the mismatch, which could be coming from e.g. background subtraction. However, as we see below, the output from GalPro is posteriors of stellar mass and redshift that are most likely broad enough to subsume the bias due to the photometry offset.
With this training set compiled, we can now train the GalPro network, keeping the default basic setting of the algorithm as described in Mucesh et al. (2020). The trained random forest is then applied to all galaxies that survive the basic cuts described in the beginning of this section. The output of each GalPro run on a single galaxy is a 2D posterior distribution of redshift and stellar mass. For each galaxy we save both the mean redshift and stellar mass values, $z_{\rm m}$ and $M_{\star,\rm m}$, which is used to define the bins as we describe in Section 4.1.2, as well as a random sample of the 2D posterior distribution, $z_{\rm samp}$ and $M_{\star,\rm samp}$. The latter, when combining a sample of galaxies, allows us to construct an estimate of the 2D distribution for the sample, as is also discussed in Section 4.1.2.

### v3

@@ 3.1 Galaxy-matter cross power spectrum $P_{\rm gm}$
Following closely our previous work in Zacharegkas2022, which follows Seljak2000; mandelbaum2004, we can write down the expressions for the terms in Equation (10) as follows.
@@
$P_{\rm gm}^{\rm 1h}(k,z)=\frac{1}{\rho_{m}n_{g}}\int dM_{h}\;\frac{dn}{dM_{h}}M_{h}$
$\quad\times u_{\rm dm}(k|M_{h})\left[\langle N_{c}(M_{h})\rangle+\langle N_{s}(M_{h})\rangle u_{\rm sat}(k|M_{h})\right]\;,$ (12)
where $u_{\rm dm/sat}(k|M_{h})$ is the Fourier transform of the dark matter/satellite spatial distribution, for which we assume the NFW profile (navarro1996), and the average galaxy number density
$n_{g}=\int dM_{h}\;\frac{dn}{dM_{h}}\langle N(M_{h})\rangle\;.$ (13)
@@
$P_{\rm gm}^{\rm 2h}(k,z)=$ $P_{\rm m}^{\rm nl}(k,z)\int dM_{h}\;\frac{dn}{dM_{h}}\frac{M_{h}}{\rho_{m}}b_{h}(M_{h})u_{\rm dm}(k|M_{h})$
$\times\int dM_{h}^{\prime}\;\frac{dn}{dM_{h}^{\prime}}\frac{b_{h}(M_{h}^{\prime})}{n_{g}}$
$\times\left[\langle N_{c}(M_{h}^{\prime})\rangle+\langle N_{s}(M_{h}^{\prime})\rangle u_{\rm sat}(k|M_{h}^{\prime})\right]\;,$ (15)
where, usually, the first and second term in the brackets are identified as the 2-halo central and 2-halo satellite terms, respectively. In this work we adopt the functions from tinker2010 for the halo bias, $b_{h}(M_{h})$, and the Halo Mass Function (HMF), $dn/dM_{h}$, from Tinker2008. We note here that on the very large scales, the above reduces to
@@
$b_{g}\equiv\frac{1}{n_{g}}\int dM_{h}\;\frac{dn}{dM_{h}}b_{h}(M_{h})\langle N(M_{h})\rangle$ (17)
is the large-scale galaxy bias, and for the linear matter power spectrum $P_{\rm m}^{\rm lin}(k,z)$ we adopt the fitting functions from eisenstein1998 which we modify by adopting the Halofit (takahashi2012) approximation to compute the nonlinear matter power spectrum, $P_{\rm m}^{\rm nl}(k,z)$. We furthermore emphasize here that the first integral in Equation (3.1) must approach unity on those linear scales. In practice, we need to choose a galaxy bias model that when combined with the halo-mass function satisfies this requirement, i.e. that the dark matter is unbiased with respect to itself, and thus:
$\lim_{k\rightarrow 0}\int dM_{h}\;\frac{dn}{dM_{h}}\frac{M_{h}}{\rho_{m}}b_{h}(M_{h})=1\;.$ (18)
@@ 3.2 Galaxy auto-correlation power spectrum $P_{\rm gg}$
To model the galaxy clustering signal, we follow vandenBosch13.
For the 1-halo components of Equation (11), we follow Seljak2000 which leads to the 1-halo galaxy auto-correlation power spectrum
$P_{\rm gg}^{\rm 1h}(k,z)=\frac{1}{n_{g}^{2}}\int dM_{h}\;\frac{dn}{dM_{h}}\lambda_{s}(M_{h})u_{\rm sat}(k|M_{h})$
$\quad\times\left[2\langle N_{c}(M_{h})\rangle+\lambda_{s}(M_{h})u_{\rm sat}(k|M_{h})\right]\;,$ (19)
@@
$\;\times\int dM_{h}\;\frac{dn}{dM_{h}}\frac{b_{h}(M_{h})}{n_{g}}\left[\langle N_{c}(M_{h})\rangle+\lambda_{s}(M_{h})u_{\rm sat}(k|M_{h})\right]$
$\;\times\int dM_{h}^{\prime}\;\frac{dn}{dM_{h}^{\prime}}\frac{b_{h}(M_{h}^{\prime})}{n_{g}}\rangle\left[\langle N_{c}(M_{h}^{\prime})\rangle+\lambda_{s}(M_{h}^{\prime})u_{\rm sat}(k|M_{h}^{\prime})\right]$
@@ 3.3 Transition regime
Following mead2021, we implement a ”beyond-linear” halo bias model when combining the 1 and 2-halo terms in Sections 3.1 and 3.2. This changes the shape of the 2-halo component as a function of $k$, especially around the scales corresponding to the size of individual dark matter halos. This scale-dependence of the halo bias becomes important on scales $r\lesssim 3\;{\rm Mpc}/h$ and we have found that it is especially important in modeling $\xi_{gg}$, as without it the 2-halo term can dominate even in the quasi-to-non-linear regime for low-mass galaxies. Moreover, this modification is especially useful for fitting our model for $\xi_{\rm gm}$ to the low-stellar mass bins, as without it our fits fail in the small-scale regime. In this paper, we implement the scale-dependent halo bias by utilizing the fitting functions from tinker2005, which are calibrated on numerical simulations and were found to be accurate enough to model galaxy clustering for galaxies in the Sloan Digital Sky Survey (SDSS; York2000_SDSS) using the HOD framework. In particular, we multiply the 2-halo correlation function by $\zeta(r,z)$, which is defined as
@@
where $\xi_{\rm m}^{\rm nl}(r,z)$ is the non-linear real-space correlation function of $P_{\rm m}^{\rm nl}(k,z)$. This is the same approach also used in zu2015 and results in the 2-halo galaxy-matter cross- and galaxy auto- correlation function terms
@@
respectively. We find that using this model we are able to obtain a better fit to our data on all scales and improve the fits significantly in the 1-halo to 2-halo regime compared to what we assumed previously in Zacharegkas2022, which was also followed by hayashi2008; zu2014; clampitt2017. As we discuss in Section 6, this model is especially useful in fitting the low-redshift, low-mass galaxy bins.
@@ 3.4 The observables: $\gamma_{t}$ and $w$
$C_{\rm gg}(\ell)=\int dz_{\ell}\;\frac{n_{\ell}^{2}(z_{\ell})}{\chi^{2}(z_{\ell})}\left(\left.\frac{d\chi}{dz}\right|_{z_{\ell}}\right)^{-1}P_{\rm gg}\left(\frac{\ell+1/2}{\chi(z_{\ell})},z_{\ell}\right)\;,$ (28)
@@ 3.5 Uncertainty in shear calibration, redshift and stellar-mass
To account for uncertainties in the shear calibration, redshift and stellar mass distributions, we make some substitutions to the equations of the previous sections, as we describe in the remaining of this section. Before we proceed, we note that aside from the marginalization of the stellar mass distribution, our approach is the same as that used in Prat2022.
@@ 3.6 Additional modeling components
Similarly to Zacharegkas2022, we consider a number of additional modeling components to our data vectors $\gamma_{t}$ and $w$. We briefly summarize each of these components and show the full model in Figure 2.
@@
where we model the 2D projected power spectrum $C_{\rm gm}^{\rm lmag}(\ell)$ following Unruh2019; Zacharegkas2022 and where $\alpha_{\rm lmag}$ is a free parameter. More specifically, in our notation, for lens bin $i$ and source bin $j$, we vary one additional parameter due to lens magnification that we call $\alpha_{\rm lmag}^{i,j}$.
Tidal stripping of satellites: As tidal forces in the outskirts of the host halo strips off the dark matter content of the satellite subhalo, the density profile of the subhalos drops off at large scales. Therefore, this term is usually modeled as a truncated NFW profile (see e.g. mandelbaum2004; velander2013) which is similar to that of the central 1-halo, $\gamma_{t}^{\rm c1h}$, out to the truncation radius $R$ and falls off as $\propto r^{-2}$ at larger radii $r$. For this work, we fix the truncation radius to $R=0.4R_{200c}$, which is roughly equivalent to having $50\%$ of the mass stripped, consistent with the average mass loss for subhaloes observed in cosmological simulations (Gao_etal_2004b; Nagai_Kravtsov_2005). Additionally, since this is a satellite term, it needs to be multiplied by the satellite fraction $F_{\rm sat}$, and we thus write:
@@
Point-mass contribution: This term quantifies the contribution to lensing by the baryonic content of the central galaxy (e.g. velander2013). This term is simply modelled as a point-source term:
$\gamma_{t}^{\rm PM}(\theta)=$ $\int dz_{\ell}\;n_{\ell}(z_{\ell})\frac{\langle M_{\star}\rangle}{\pi r^{2}(\theta;z_{\ell})}\int dz_{s}n_{s}(z_{s})\Sigma_{c}^{-1}(z_{\ell},z_{s})\;.$ (38)
In this work, $\langle M_{\star}\rangle$ is the average stellar mass content in a given lens bin from Table 1. The final contribution of this term to $\gamma_{t}$ is computed as in Zacharegkas2022. There is no free parameters in this component.
Density profiles for satellite galaxies: For the spatial distribution of dark matter and satellite galaxies within a dark matter halo we have assumed the NFW proflie (navarro1996). Thus, their Fourier transforms have analytic forms with concentration parameters $c_{\rm dm}$ and $c_{\rm sat}$, respectively for dark matter and satellite galaxies. However, the distribution of satellite galaxies is typically less concentrated than that of the dark matter (e.g. Carlberg1997; Lin2004; Nagai2004; Hansen2004). To account for this, in this work we allow $c_{\rm sat}$ to be different than $c_{\rm dm}$ by introducing the free parameter $\kappa_{c}^{i}=c_{\rm sat}/c_{\rm dm}$ per lens bin $i$.
Intrinsic alignment: Galaxies are not randomly oriented even in the absence of lensing. On large scales, galaxies can be stretched in a preferable direction by the tidal field of the large scale structure. On small scales, other effects such as the radial orbit of a galaxy in a cluster can affect their orientation. This phenomenon, where the shape of the galaxies is correlated with the density field, is known as intrinsic alignment (IA, troxel2015). The contamination of shear by IA can become important when the source galaxies are physically close to the lenses and gravitational interactions can modify the shape of the galaxies. A good proxy for whether IA’s contribution to $\gamma_{t}$ is significant is the measured boost factors (see Section 5.4 and Appendix E for the relevant discussions and definitions). Therefore, given our bin selection and scale cuts, which are described in Section 5.4, we have ignored the contribution from IA in this work.
4 Data
In this section we introduce the data used in this work. This work is based on the first three years (Y3) of DES data (DESY3GOLD) and inherits all the infrastructure and value-added products associated with that dataset (e.g. DESY3shear; Myles2021; DESY3DeepFields). The most novel aspect in the dataset here is the new stellar mass-selected lens sample we construct and validate in this work. The source sample follows the same selection as described in Zacharegkas2022; Prat2022. As such, we describe in more detail the lens sample in Section 4.1 and briefly introduce the source sample in Section 4.2.
4.1 The stellar mass-selected lens sample
To generate our stellar mass-selected sample, we start from the DES Y3 GOLD catalog (DESY3GOLD). We first define an approximately complete galaxy sample by performing the following steps:
  • Select all objects with i-band magnitude sof_cm_mag_corrected_i<23<23
  • Define footprint to only contain regions with limiting magnitude deeper than sof_cm_mag_corrected_i<23<23
@@ 4.1 The stellar mass-selected lens sample
This catalog is then fed into the GalPro (Mucesh2020) package to generate photometric redshift and stellar mass for each galaxy. We briefly introduce GalPro and our training set in Section 4.1.1, then define the actual sample used for this work in Section 4.1.2. In Section 4.1.3 we describe the procedure in which we obtain weights for this lens sample to ensure it is not significantly impacted by the spatial variation of observing conditions.
@@ 4.1.1 GalPro and the training set
GalPro (Mucesh2020) is a python-based machine learning code based on the Random Forest algorithm for estimating multivariate posterior probability distributions of galaxy properties. It can operate generically on a wide range of galaxy properties (e.g. redshift, stellar mass, star formation rate, metallicity), while we use it to extract redshift and stellar mass information.
The input quantities that GalPro accepts are luptitudes (Lupton1999), $\mu$, their uncertainties $\sigma_{\mu}$, as well as lupti-colors and their uncertainties. Luptitudes behave like magnitudes for bright photometry and like fluxes for faint photometry, with the turning point in the behaviour determined by the softening parameter. Thus, by converting to luptitudes, we avoid introducing an additional selection effect by not discarding galaxies with negative fluxes.
@@
For the colors, we calculate three lupti-colors grg-r, rir-i and izi-z and their corresponding errors. These are simply calculated as $\mu_{\rho}-\mu_{\nu}$ for two bands $\rho,\nu\in\{g,r,i,z\}$ and $\sqrt{\sigma_{\mu,\rho}^{2}+\sigma_{\mu,\nu}^{2}}$, respectively.
Before we can apply GalPro to the DES data, we need to train the random forest with a training set where we have “truth” information for the mapping between the photometry and the desired quantities (redshift and stellar mass in our case). We choose to train our model on galaxies in the DES Y3 Deep Fields (DF, DESY3DeepFields), matched to an overlapping subset of galaxies in the Cosmic Evolution Survey COSMOS2020 (COSMOS, COSMOS2020). In matching the two samples, we adjust the stellar mass slightly according to the difference in amplitude between the DF photometry and the COSMOS photometry. We note that this may not address the underlying cause of the mismatch, which could be coming from e.g. background subtraction. However, as we see below, the output from GalPro is posteriors of stellar mass and redshift that are most likely broad enough to subsume the bias due to the photometry offset.
With this training set compiled, we can now train the GalPro network, keeping the default basic setting of the algorithm as described in Mucesh2020. The trained random forest is then applied to all galaxies that survive the basic cuts described in the beginning of this section. The output of each GalPro run on a single galaxy is a 2D posterior distribution of redshift and stellar mass. For each galaxy we save both the mean redshift and stellar mass values, $z_{\rm m}$ and $M_{\star,\rm m}$, which is used to define the bins as we describe in Section 4.1.2, as well as a random sample of the 2D posterior distribution, $z_{\rm samp}$ and $M_{\star,\rm samp}$. The latter, when combining a sample of galaxies, allows us to construct an estimate of the 2D distribution for the sample, as is also discussed in Section 4.1.2.